OpenStudy (anonymous):

Create an absolute value equation that has one solution. Explain why this equation has one solution.

OpenStudy (dumbcow):

|x+5| = 0 x = -5 only has 1 solution because +0 is same as -0

OpenStudy (anonymous):

Thanks .. : )

hero (hero):

What? 0 is neither positive or negative. Dumbcow I'm surprised at you.

OpenStudy (dumbcow):

lol :) ok sorry i just was relating to normally solving these you would have a pos and neg case yes zero is neither pos or neg, hence only 1 possible case

OpenStudy (anonymous):

I understood.

OpenStudy (dumbcow):

graphically for it to have 1 solution the vertex would have to be resting on the x-axis, thus the y_value would have to be 0

OpenStudy (anonymous):

@Hero: What? +0 = -0 is a statement about zero and the unary minus ... compare to 8+0 = 8-0 ... and makes no claims about "positive" or "negative" ... -5 = +5 is false while -0 = +0 is true